# Stitched ViTs are Flexible Vision Backbones

### v1

@@ Abstract
Large pretrained plain vision Transformers (ViTs) have been the workhorse for many downstream tasks. However, existing works utilizing off-the-shelf ViTs are inefficient in terms of training and deployment, because adopting ViTs with individual sizes requires separate training and is restricted by fixed performance-efficiency trade-offs. In this paper, we are inspired by stitchable neural networks, which is a new framework that cheaply produces a single model that covers rich subnetworks by stitching pretrained model families, supporting diverse performance-efficiency trade-offs at runtime. Building upon this foundation, we introduce SN-Netv2, a systematically improved model stitching framework to facilitate downstream task adaptation. Specifically, we first propose a Two-way stitching scheme to enlarge the stitching space. We then design a resource-constrained sampling strategy that takes into account the underlying FLOPs distributions in the space for improved sampling. Finally, we observe that learning stitching layers is a low-rank update, which plays an essential role on downstream tasks to stabilize training and ensure a good Pareto frontier. With extensive experiments on ImageNet-1K, ADE20K, COCO-Stuff-10K, NYUv2 and COCO-2017, SN-Netv2 demonstrates strong ability to serve as a flexible vision backbone, achieving great advantages in both training efficiency and adaptation. Code will be released at https://github.com/ziplab/SN-Netv2.
@@ 1 Introduction
General-purpose Transformer architectures [3, 15, 12, 53, 26, 5] have grown into unprecedented scale in recent research. In computer vision, large pretrained plain ViTs such as MAE [20], DINO [5, 32] and DeiT [44, 45] are widely adopted as backbones for tackling downstream tasks. However, despite the great performance, when adopting pretrained ViTs into the downstream tasks, they all face the challenge of the huge computational and memory cost. For example, processing a single $224\times 224$ image using DeiT3-Large [45] requires 2.8G peak GPU memory consumption, which is 14$\times$ higher than that of using DeiT3-Small (0.2G).
@@
On the other hand, most existing efforts adopting Transformers as backbones for downstream tasks can be roughly categorised into three types of approaches: full-adoption [28, 56, 42, 16], parameter-efficient tuning [24, 7] and adapters [9]. Specifically, full-adoption trains one scale of ViT on downstream tasks while parameter-efficient methods additionally reduce the number of learnable parameters. Adapter method like ViT-Adapter [9] equips plain ViTs with pyramid feature maps for dense predictions. However, these methods do not address the memory consumption and computational cost after adoption.
@@
Most recently, a promising framework has been proposed named stitchable neural networks (SN-Netv1, Figure 2 (a)) [33], where it reassembles a new network (stitch) by stitching pretrained model families (anchors) in a Fast-to-Slow direction, i.e., the stitch begins with a small ViT and goes into a large ViT via a stitching layer ($1\times 1$ convolution). At training time, SN-Net randomly samples a stitch and train it as a normal network with gradient descent. With a few epochs of finetuning on the same pretraining dataset (i.e., ImageNet-1K), SN-Netv1 produces numerous networks with different performance-efficiency trade-offs that satisfy various resource constraints. However, many general-purpose ViTs are trained on exceptionally large datasets (e.g., JFT [53], LAION [39]) by industry, making it impractical for researchers to train SN-Netv1 on the same domain as the anchors. Therefore, exploring the downstream adaptation of SN-Net becomes essential as it enables researchers to use off-the-shelf pretrained models without significant retraining overhead on the original dataset.
In this paper, we systematically improve SN-Netv1 for direct adaptation on downstream dense predictions. The new method, called SN-Netv2, presents impressive power as a strong and flexible vision backbone on downstream CV tasks. Specifically, our improvements are in three folds: 1) Different from SN-Netv1 that only permits the stitching direction of Fast-to-Slow, we propose a novel Two-way stitching strategy (Figure 1), which enables stitching to go both Fast-to-Slow and Slow-to-Fast, or travel between anchors as a round trip (e.g., Fast-Slow-Fast). As a result, we effectively enlarge the stitching space by 10×10$\times$10 × and avoid sub-optimal stitches that reside in certain resource constraints. However, under the default random sampling strategy, simply enlarging the space incurs unbalanced training and hinders the overall performance of SN-Netv1 due to the varying number of stitches residing on different FLOPs intervals. To this end, 2) we introduce a Resource-constrained Sampling (ROS) strategy, which draws a stitch at each iteration according to the categorical distribution of stitches in the space [46, 30]. In this case, ROS ensures a balanced training for stitches under different resource constraints. 3) Last, although SN-Netv1 demonstrated the advantage of fully finetuning the stitching layers after least-square initialization, we point out that learning stitching layers resides on a low intrinsic dimension. On ImageNet-1K, low-rank adaptation of stitching layers achieves competitive performance, while it plays a key role in stabilizing training and ensuring a smooth performance curve on downstream tasks.
With the improved techniques, we comprehensively experiment with SN-Netv2 on ImageNet-1K, ADE20K, COCO-Stuff-10K, NYUv2 and COCO. We show SN-Netv2 demonstrates strong performance and training efficiency than typical single-scale ViT backbone adoption on downstream dense prediction tasks. In particular, on ImageNet-1K, SN-Netv2 obtains better performance-efficiency trade-offs than SN-Netv1. On downstream tasks, SN-Netv2 achieves competitive performance with anchors under equal training schedules, while eliminating the need to train different scales of ViT backbone separately. With systematic evaluation, we show SN-Netv2 enables flexible inference on memory and computation-intensive dense predictions with a single network, showcasing its potential as a next-generation backbone for a wide range of tasks.
2 Related Work
General-purpose Transformers.
Benefit from the large-scale datasets [39, 53] and powerful scaling laws [25], recent pretrained plain ViTs [20, 2, 14, 12] have achieved strong performance on many visual benchmarks. Based on different training objectives, ViT pretraining can be done by either supervised learning (SL) or self-supervised learning (SSL). In SL-based pretraining, the common practice [31, 14] is to train ViTs with the cross-entropy loss on ImageNet-21K [38] or privately gathered images [43]. Compared to it, SSL-based pretraining is a more promising direction as it can leverage the vast amounts of unlabeled data. Prevalent approaches in this area include contrastive learning [5, 8, 52, 35] and masked image modeling (MIM) [20, 2, 58, 13, 18, 50]. In recent studies, ViTs have been scaled up to billion parameters [53, 12, 15] to match the prevalent large language models (LLMs) [55, 36, 3]. However, when adopting into downstream tasks, most existing efforts [28, 16, 54, 56] fully adopt the pretrained model in order to exploit the pretrained weights, suffering from the huge computational cost and memory consumption. In contrast, this paper efficiently adapts large pretrained ViTs into CV tasks as a single flexible backbone to satisfy various resource constraints at runtime.
Model stitching.
Model stitching has been studied in previous works [27, 1, 11] to measure the similarity of representations from neural networks. In particular, it implies a sequence of well-performed networks that can be obtained by stitching the early portion of a trained network with the last portion of another trained network by a $1\times 1$ convolution layer (a.k.a, stitching layer). Inspired by this observation, Yang et al.  [51] proposed to dissect and reassemble a new architecture based on the model zoo. Most recently, Pan et al. proposed SN-Net [33] to cheaply produce numerous networks with different complexity-performance trade-offs by stitching a family of pretrained models. However, it only experiments with the same pretraining domain, without exploring dense prediction performance. In this paper, we systematically improves SN-Net and address its limitations on the stitching space, sampling strategy and downstream training adaptation.
3 Rethinking Stitchable Neural Networks
Figure 2: Explanation of SN-Netv1 [33]. Figure (a): The framework of SN-Netv1 under the default Fast-to-Slow strategy, where the blue line indicates the forward path of a stitched network by stitching a small ViT to a large ViT. Figure (b): The naive Fast-to-Slow stitching strategy incurs some bad stitches (highlight in red box) when stitching pretrained DeiT3-S and DeiT3-L on ImageNet-1K.
@@ 3 Rethinking Stitchable Neural Networks
Stitching layers. SN-Netv1 is initially trained on the same data source of the pretrained model families (e.g., ImageNet-1K pretrained DeiTs [44]). Under this setting, training a stitching layer with fully finetuning has a consistent target as the activation space $\mathcal{A}_{\theta,l}$ and $\mathcal{A}_{\phi,m}$ may not change significantly. However, things can be different when adopting SN-Netv1 on downstream dense prediction tasks due to the domain gap between the pretrained data and the target data. In this case, both $\mathcal{A}_{\theta,l}$ and $\mathcal{A}_{\phi,m}$ need to adapt to the target domain, making it unstable and difficult to simultaneously learn many stitches. This implies the necessity of an appropriate method for learning stitching layers on downstream tasks.
@@ 4.1 Two-way Stitching
To improve the stitching space, we first propose Two-way stitching, which allows the stitching to travel in different ways including Fast-to-Slow (FS), Slow-to-Fast (SF), Fast-Slow-Fast (FSF) and Slow-Fast-Slow (SFS). The design principle of Two-way Stitching is to augment the small anchor with the blocks from a large and strong anchor, while accelerating the large anchor by adopting more efficient blocks from the small anchor. In this way, Two-way stitching can leverage the strengths of different scales of anchors.
We illustrate our framework in Figure 1, where it also shows a concrete example of an FSF stitch in the blue line of Figure 1. Specifically, its begins with the small anchor during the forwarding pass, traverses the large anchor along the middle route, and ultimately returns to the small anchor for subsequent propagations. By enabling a broader stitching configuration, Two-way stitching enlarges the stitching space by 10×10$\times$10 × compared to the initial space of SN-Netv1 (e.g., 134 vs. 13 based on DeiT3-S/L), which facilitates the discovery of more optimal architectures, as shown in Section 5. Moreover, benefiting from the two new stitching configurations (i.e., FSF and SFS), SN-Netv2 can produce many networks at the similar FLOPs as each stitch can be regarded as replacing intermediate consecutive blocks at certain positions in one anchor with the blocks from another anchor.
@@ 4.2 Resource-Constrained Sampling
After enlarging the stitching space, we empirically find that the performance of anchors drops significantly compared to its original performance. To understand this, we begin with analysing the categorical distribution of stitches in the space, i.e., $\pi(\tau)$, where $\tau$ denotes the FLOPs constraint. In practice, we round the real FLOPs following a step $t$ to discretize the whole FLOPs range. By default, we adopt $t=1$ in ImageNet experiments and $t=10$ for dense prediction tasks. Benefiting from the significantly smaller architectural space compared to neural architecture search (i.e., NAS, $10^{2}$ vs. $10^{20}$), we can calculate $\pi(\tau)$ exactly prior to training by $\pi(\tau=\tau_{0})=\#(\tau=\tau_{0})/E$, where $E$ is the total number of stitches in the space and #$\#(\tau=\tau_{0})$ is the number of stitches that yield FLOPs $\tau_{0}$.
@@ 4.3 Low-Rank Adaptation of Stitching Layers
However, as mentioned in Section 3, learning a good stitching layer on downstream tasks can be difficult as anchors need to adapt to the target domain simultaneously. Motivated by previous observations [17, 6] that sparse finetuning via paremeter-efficient tuning (PET) can impose regularization on the model which facilitates training stability, we propose LoRA SL, a low-rank adaptation method for stitching layers in order to stabilize training. Specifically, similar to LoRA [23], we freeze the LS initialized transformation matrix $\mathbf{M}$ but constrain its update with a low-rank decomposition. Taking $\mathcal{S}_{1}$ as a concrete example, the activation projection of the stitching layer can be formulated by
@@
where $\mathbf{B}_{1}\in\mathbb{R}^{D_{1}\times r}$, $\mathbf{A}_{1}\in\mathbb{R}^{r\times D_{2}}$, and $r\ll min(D_{1},D_{2})$ is the rank. In practice, $\mathbf{B}_{1}$ is initialized by Gaussian initialization and $\mathbf{A}_{1}$ is initialized with zeros. Therefore, the initial update $\Delta\mathbf{M}_{1}=\mathbf{B}_{1}\mathbf{A}_{1}$ is zero and does not affect the LS solution at the beginning of training. As each stitiching layer is responsible for multiple stitches, the low-rank update helps to regularize the learning of $\mathbf{M}_{1}$. We show in Figure 11 that LoRA SL improves the overall performance of SN-Netv2. In Algorithm 1, we summarize our training approach for SN-Netv2 and highlight the difference with SN-Netv1 in bold. It is also worth noting that we do not adopt knowledge distillation to train SN-Netv2 as it is very inefficient on downstream dense prediction tasks.
@@
Figure 5: Semantic segmentation results of SN-Netv2 on ADE20K by stitching DeiT3-S/L and DeiT3-B/L. We use yellow stars to denote the anchor performance, i.e., adopt the individual anchor as a backbone for training. We highlight the stitches at the Pareto frontier in blue line.

### v2

@@ Abstract
Large pretrained plain vision Transformers (ViTs) have been the workhorse for many downstream tasks. However, existing works utilizing off-the-shelf ViTs are inefficient in terms of training and deployment, because adopting ViTs with individual sizes requires separate trainings and is restricted by fixed performance-efficiency trade-offs. In this paper, we are inspired by stitchable neural networks (SN-Net), which is a new framework that cheaply produces a single model that covers rich subnetworks by stitching pretrained model families, supporting diverse performance-efficiency trade-offs at runtime. Building upon this foundation, we introduce SN-Netv2, a systematically improved model stitching framework to facilitate downstream task adaptation. Specifically, we first propose a two-way stitching scheme to enlarge the stitching space. We then design a resource-constrained sampling strategy that takes into account the underlying FLOPs distributions in the space for better sampling. Finally, we observe that learning stitching layers as a low-rank update plays an essential role on downstream tasks to stabilize training and ensure a good Pareto frontier. With extensive experiments on ImageNet-1K, ADE20K, COCO-Stuff-10K and NYUv2, SN-Netv2 demonstrates superior performance over SN-Netv1 on downstream dense predictions and shows strong ability as a flexible vision backbone, achieving great advantages in both training efficiency and deployment flexibility. Code is available at https://github.com/ziplab/SN-Netv2.
Figure 1: We adapt the framework of stitchable neural networks (SN-Net) into downstream dense prediction tasks. Compared to SN-Netv1 [33], the new framework consistently improves the performance at low FLOPs while maintaining competitive performance at high FLOPs across different datasets, thus obtaining a better Pareto frontier (highlighted in lines). Compared to individual large ViT trainings, SN-Netv2 obtains numerous trade-offs with less A100 GPU training hours. At the inference time, SN-Netv2 inherits the advantage of SN-Net which allows runtime speed-performance trade-off switching to satisfy different resource constraints. Besides ImageNet-1K classification, we evaluate the performance of semantic segmentation in mIoU on ADE20K and COCO-Stuff-10K, and depth estimation in RMSE on NYUv2.
[…]
@@ 1 Introduction
General-purpose Transformer architectures [3, 15, 12, 53, 26, 5] have grown into unprecedented scale in recent research. In computer vision, large pretrained plain ViTs such as MAE [21], DINO [5, 32] and DeiT [44, 45] are widely adopted as backbones for tackling downstream tasks. However, despite the great performance, when adapting large pretrained ViTs to the downstream tasks, they face the challenge of huge computational cost. For example, for semantic segmentation on COCO-Stuff-10K, DeiT3-Large [45] suffers from longer fine-tuning time (90 GPU hours) and slow inference speed (5.7 FPS), compared to DeiT3-S (40 GPU hours and 28.2 FPS), see Fig. 1.
@@
On the other hand, most existing efforts adopting pretrained ViTs as backbones for downstream tasks can be roughly categorised into three types of approaches: full finetuning [28, 56, 42, 16], parameter-efficient finetuning [24, 7] and adapters [9]. Specifically, full finetuning trains one scale of ViTs on downstream tasks while parameter-efficient methods additionally reduce the number of learnable parameters. Adapter methods like ViT-Adapter [9] equip plain ViTs with pyramid feature maps for dense predictions. However, these methods lack deployment flexibility since they cannot cope with diverse deployments with different computing budget requirements.
Figure 2: Comparison between SN-Netv1 and SN-Netv2. Different from SN-Netv1 which only considers the classification task on the same data domain of the pretrained model family, SN-Netv2 solves the limitations of SN-Netv1 in downstream tasks and successfully extends the practical usage of the framework.
Recently, a flexible deployment framework named stitchable neural network (SN-Netv1 [33]) has been proposed, which leverages a pretrained model family to obtain numerous sub-networks with different efficiency-performance trade-offs. A simple illustration is shown in Fig. 3 (a), where SN-Netv1 inserts a few stitching layers (1 $\times$ 1 convolutional layers) between two scales of pretrained ViTs (i.e., “anchors” in SN-Netv1). Next, SN-Netv1 adopts a Fast-to-Slow strategy which forwards the activations from the small ViT into the large ViT via different stitching layers, thus obtaining many sub-networks (i.e., “stitches”). With a few epochs of finetuning on the same pretraining dataset as anchors, SN-Netv1 can instantly switch network topology at runtime under various resource constraints.
Despite offering a flexible deployment framework, SN-Netv1 suffers from several major limitations. First, SN-Netv1 performs full finetuning on the same original dataset as the anchors. This might not be practical since many general-purpose ViTs are trained on exceptionally large datasets (e.g., JFT [53], LAION [39]). Second, SN-Netv1 only investigates image classification, whose data domain is different from downstream task data domains. Third, due to the small stitching space, SN-Netv1 can produce sub-optimal performance-efficiency trade-offs, and the domain gap can lead to more severe situations on downstream dense predictions, as shown in Fig. 3 (b) and (c).
In this paper, we introduce a new framework called SN-Netv2, serving as a strong and flexible vision backbone on downstream vision tasks, which systematically improves SN-Netv1 for direct adaptation on downstream dense predictions to obtain a smooth and high-performance curve across all FLOPs (see Fig. 1 and Fig. 2). Specifically, we propose three tightly coupled improvements: 1) Different from SN-Netv1 that only permits the stitching direction of Fast-to-Slow, we propose a novel Two-way Stitching (TWS) strategy (Fig. 4), which enables stitching to go both Fast-to-Slow and Slow-to-Fast, or travel between anchors as a round trip (e.g., Fast-Slow-Fast). As a result, we effectively enlarge the stitching space by 10×10$\times$10 × and avoid sub-optimal stitches that reside in certain resource constraints. However, under the default random sampling strategy, simply enlarging the space incurs unbalanced training and hinders the overall performance due to the varying number of stitches residing on different FLOPs intervals. 2) To this end, we further introduce a Resource-constrained Sampling (ROS) strategy, which draws a stitch at each iteration according to the categorical distribution of stitches in the space [46, 29] so as to ensure a balanced training for stitches under different resource constraints. 3) Last, to address the training instability caused by the domain gap between anchors and the target task, along with the increased interference among stitches in the enlarged stitching space, unlike the full-finetune strategy of SN-Netv1, we propose to stabilize training with a low-rank adaption of stitching layers (LoRA SL), which ensures a smooth performance curve. Note that we utilize LoRA as a novel solution for the critical training challenge of adapting SN-Net to new downstream tasks, which is different from exploring LoRA [23] for parameter-efficient finetuning in existing methods [7, 31].
Figure 3: Explanation of SN-Netv1 [33]. Figure (a): The framework of SN-Netv1 under the default Fast-to-Slow strategy, where the blue line indicates the forward path of a stitched network by stitching a small ViT to a large ViT. Figure (b): The naive Fast-to-Slow stitching strategy results in sub-optimal trade-offs in the low-FLOPs range on the pretrained dataset. Figure (c): Directly applying SN-Netv1 into ADE20K obtains a bad frontier, especially when comparing with the individually trained small ViT (40.4 vs. 45.5 mIoU).
With the improved techniques, we comprehensively experiment with SN-Netv2 on ImageNet-1K, ADE20K, COCO-Stuff-10K, NYUv2 and COCO-2017. We show SN-Netv2 demonstrates stronger performance and better training efficiency than the typical single-scale ViT backbone adoption on downstream dense prediction tasks. Particularly, on both the pretraining and downstream tasks, SN-Netv2 consistently outperforms SN-Netv1 at low FLOPs, while keeping competitive performance at high FLOPs. Moreover, on downstream dense predictions, SN-Netv2 achieves competitive performance with anchors under equal training schedules, while eliminating the need to train different scales of ViT backbones separately.
Our main contributions can be summarized as follows:
  • Compared to SN-Netv1 which only considers the classification task on the same data domain as pretrained models, we explore the applicability of SN-Net in downstream dense predictions.
  • To overcome the suboptimal trade-offs inherent in SN-Netv1 in downstream tasks, we propose SN-Netv2, which seamlessly integrates three tightly coupled components, TWS, ROS and LORA SL, and achieves a superior Pareto frontier in terms of performance and efficiency.
  • We conduct comprehensive experiments across various datasets and tasks to evaluate SN-Netv2, showcasing its great potential as a strong and flexible vision backbone for downstream vision tasks.
2 Related Work
General-purpose Transformers. Benefit from the large-scale datasets [39, 53] and powerful scaling laws [25], recent pretrained plain ViTs [21, 2, 14, 12] have achieved strong performance on many visual benchmarks. Based on different training objectives, ViT pretraining can be done by either supervised learning (SL) or self-supervised learning (SSL). In SL-based pretraining, the common practice [30, 14] is to train ViTs with the cross-entropy loss on ImageNet-21K [38] or privately gathered images [43]. Compared to it, SSL-based pretraining is a more promising direction as it can leverage the vast amounts of unlabeled data. Prevalent approaches in this area include contrastive learning [5, 8, 52, 36] and masked image modeling (MIM) [21, 2, 58, 13, 18, 50]. In recent studies, ViTs have been scaled up to billion parameters [53, 12, 15] to match the prevalent large language models (LLMs) [55, 35, 3]. However, when adopting into downstream tasks, most existing efforts [28, 16, 54, 56] fully adopt the pretrained model in order to exploit the pretrained weights, suffering from the huge computational cost and memory consumption. In contrast, this paper efficiently adapts large pretrained ViTs into CV tasks as a single flexible backbone to satisfy various resource constraints at runtime.
Model stitching. Model stitching has been studied in previous works [27, 1, 11] to measure the similarity of representations from neural networks. In particular, it implies a sequence of well-performed networks that can be obtained by stitching the early portion of a trained network with the last portion of another trained network by a $1\times 1$ convolution layer (a.k.a, stitching layer). Inspired by this observation, Yang et al. [51] proposed to dissect and reassemble a new architecture based on the model zoo. Most recently, Pan et al. proposed SN-Net [33] to cheaply produce numerous networks with different complexity-performance trade-offs by stitching a family of pretrained models. However, it only experiments with the same pretraining domain, without exploring dense prediction performance. In this paper, we systematically improves SN-Net and address its limitations on the stitching space, sampling strategy and downstream training adaptation.
3 Rethinking Stitchable Neural Networks
@@
Stitching layers. SN-Netv1 is initially trained on the same data source of the pretrained model families (e.g., ImageNet-1K pretrained DeiTs [44]). Under this setting, training a stitching layer with full finetune has a consistent target as the activation space $\mathcal{A}_{\theta,l}$ and $\mathcal{A}_{\phi,m}$ may not change significantly. However, things can be different when adopting SN-Netv1 on downstream dense prediction tasks due to the domain gap between the pretrained data and the target data. In this case, both $\mathcal{A}_{\theta,l}$ and $\mathcal{A}_{\phi,m}$ need to adapt to the target domain, making it unstable and difficult to simultaneously learn many stitches. This implies the necessity of an appropriate method for learning stitching layers on downstream tasks.
@@ 4.1 Two-way Stitching
To improve the stitching space, we first propose Two-way stitching (TWS), which allows the stitching to travel in different ways including Fast-to-Slow (FS), Slow-to-Fast (SF), Fast-Slow-Fast (FSF) and Slow-Fast-Slow (SFS). The design principle of TWS is to augment the small anchor with the blocks from a large and strong anchor, while accelerating the large anchor by adopting more efficient blocks from the small anchor. In this way, TWS can leverage the strengths of different scales of anchors.
We illustrate our framework in Figure 4, where it also shows a concrete example of an FSF stitch in the blue line of Figure 4. Specifically, its begins with the small anchor during the forwarding pass, traverses the large anchor along the middle route, and ultimately returns to the small anchor for subsequent propagations. By enabling a broader stitching configuration, TWS enlarges the stitching space by 10×10$\times$10 × compared to the initial space of SN-Netv1 (e.g., 134 vs. 13 based on DeiT3-S/L), which facilitates the discovery of more optimal architectures, as shown in Section 5. Moreover, benefiting from the two new stitching configurations (i.e., FSF and SFS), SN-Netv2 produces many networks at the similar FLOPs as each stitch can be regarded as replacing intermediate consecutive blocks at certain positions in one anchor with the blocks from another anchor.
@@ 4.2 Resource-Constrained Sampling
After enlarging the stitching space, we empirically find that the performance of anchors drops significantly compared to its original performance on ImageNet. To understand this, we begin with analysing the categorical distribution of stitches in the space, i.e., $\pi(\tau)$, where $\tau$ denotes the FLOPs constraint. In practice, we round the real FLOPs following a step $t$ to discretize the whole FLOPs range. By default, we adopt $t=1$ in ImageNet experiments and $t=10$ for dense prediction tasks. Benefiting from the significantly smaller architectural space compared to neural architecture search (i.e., NAS, $10^{2}$ vs. $10^{20}$), we can calculate $\pi(\tau)$ exactly prior to training by $\pi(\tau=\tau_{0})=\#(\tau=\tau_{0})/E$, where $E$ is the total number of stitches in the space and #$\#(\tau=\tau_{0})$ is the number of stitches that yield FLOPs $\tau_{0}$.
@@ 4.3 Low-Rank Adaptation of Stitching Layers
However, as mentioned in Section 3, learning a good stitching layer on downstream tasks can be difficult as anchors need to adapt to the target domain simultaneously. Motivated by previous observations [17, 6] that sparse finetuning via paremeter-efficient finetuning (PEFT) can impose regularization on the model which facilitates training stability, we propose LoRA SL, a low-rank adaptation method for stitching layers in order to stabilize training. Specifically, similar to LoRA [23], we freeze the LS initialized transformation matrix $\mathbf{M}$ but constrain its update with a low-rank decomposition. Taking $\mathcal{S}_{1}$ as a concrete example, the activation projection of the stitching layer can be formulated by
@@
where $\mathbf{B}_{1}\in\mathbb{R}^{D_{1}\times r}$, $\mathbf{A}_{1}\in\mathbb{R}^{r\times D_{2}}$, and $r\ll min(D_{1},D_{2})$ is the rank. In practice, $\mathbf{B}_{1}$ is initialized by Gaussian initialization and $\mathbf{A}_{1}$ is initialized with zeros. Therefore, the initial update $\Delta\mathbf{M}_{1}=\mathbf{B}_{1}\mathbf{A}_{1}$ is zero and does not affect the LS solution at the beginning of training. As each stitiching layer is responsible for multiple stitches, the low-rank update helps to stabilize training. We show in Figure 10 that LoRA SL improves the overall performance of SN-Netv2. In Algorithm 1, we summarize our training approach for SN-Netv2 and highlight the difference with SN-Netv1 in bold. It is also worth noting that we do not adopt knowledge distillation to train SN-Netv2 as it is very inefficient on downstream dense prediction tasks.
@@
Figure 7: Semantic segmentation results of SN-Netv1 and SN-Netv2 on ADE20K by stitching DeiT3-S/L and DeiT3-B/L. We highlight the stitches at the Pareto frontier in lines.
Figure 7: Semantic segmentation results of SN-Netv1 and SN-Netv2 on ADE20K by stitching DeiT3-S/L and DeiT3-B/L. We highlight the stitches at the Pareto frontier in lines.
Figure 8: Semantic segmentation results of SN-Netv1 and SN-Netv2 on COCO-Stuff-10K by stitching DeiT3-S/L and DeiT3-B/L. We highlight the stitches at the Pareto frontier in lines.
Figure 9: Results of stitching DeiT3-S and DeiT3-L on NYUv2 [40] under the framework of DPT [37].